# Maximum Inner Product is Query-Scaled Nearest Neighbor

### v1

@@ Abstract.
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
(2) Scalable Solution: We introduce $\mathsf{PSP}$, a scalable graph index with spherical pathways and adaptive search initialization for the biased distribution of $\mathsf{MIPS}$ solutions. We propose an adaptive early termination mechanism to mitigate the excessive exploration. (3) Extensive Experiments on 8 real-world datasets validate PSP’s theoretical soundness, efficiency, and scalability, highlighting an average 35% speed-up in query process and a 3$\times$ reduction in index size over leading graph methods. PSP is the only one reaching 95% recall under 5ms and thus has been deployed in the large-scale search engines of Shopee, which showcases the application viability.
@@ 2. Preliminaries and Background
While defining $k$-$\mathsf{MIPS}$ and $k$-$\epsilon$$\mathsf{MIPS}$ to find the top $k$ solutions is straightforward, we omit these details here for brevity. Note that the main difference between MIPS and a similar field, Nearest Neighbor Search (NNS), lies in the metric used (Pan et al., 2023) (inner product and Euclidean distance respectively) [^1] MIPS is often used in recommender and search engine, where relevance is key (Huang et al., 2020). NNS emphasizes spatial proximity and is used in classification and clustering (Philbin et al., 2007).. Recent efforts aim to address both $\epsilon$NNS and $\epsilon$$\mathsf{MIPS}$ problems with similar strategies: reducing the cost of distance calculations and pruning the search space (Zhao et al., 2023a; Ghojogh et al., 2018; Guo et al., 2020; Zheng et al., 2016). To better motivate our study, we revisit the $\epsilon$$\mathsf{MIPS}$ techniques from a novel view: whether transforming the vector space.
@@ 3.1. Equivalence under Scaled Query
We exclude cases where the zero-vector or multiple identical vectors in $\mathcal{D}$ complicate the analysis. The IP of the zero vector with any vector equals 00, which lacks practical meaning.
Then, we break down the proof step by step.
@@
To align the search paths under both metrics as per Algorithm 1, we can unify the node-selection behavior by localizing the problem: solving for the optimal $p^{*}$ among the neighbors at each greedy step, akin to the conditions established in Theorem 1. Let $l$ be the number of steps in the path. Define $U=\left\{\bar{\mu}_{i}|1\leq i\leq l\right\}$. it’s not hard to verify that there exists a scalar boundary $\sup\left(U\cup\{0\}\right)$ forces the Algorithm 1 under both IP and Euclidean metrics to generate the same search path targeting query $\mu q$. Please refer to appendix for the complete proof. ∎
Remarks. From Theorem 1 and Theorem 2, we discern a compelling duality between the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paradigms. These findings suggest that by appropriately scaling the query $q$, we can modulate the overlap in their search paths on a Euclidean proximity graph. By choosing an appropriate $\mu$, the search behavior becomes invariant to the metric used. Our experimental validations further corroborate this theoretical insight on real-world datasets (§5.2).
@@ 3.2. Efficiency Analyses for Graph-Based MIPS
The complete process of derivation can be found in appendix. ∎
@@ 7 $G\leftarrow\{G,N\}$
Remarks. Previous transformation-based methods (Shrivastava and Li, 2014, 2015; Neyshabur and Srebro, 2015; Li et al., 2018; Yan et al., 2018; Zhao et al., 2023a; Zhou et al., 2019) sought to reduce the $\mathsf{MIPS}$ problem to the $\mathsf{NNS}$ problem using non-linear transformations, suffering from strong theoretical assumptions (Zhou et al., 2019), structural distortions (Zhao et al., 2023a; Morozov and Babenko, 2018), and complicate data updates (§2), while failed to provide theoretical guarantees on the reachability of the 1-$\mathsf{MIPS}$ solution. Our Theorems 1 and 2 demonstrate that this reachability can be achieved without space transformations. Theorems 3 and 4 validate the efficiency of solving $\mathsf{MIPS}$ on an SSG, highlighting the crucial role of edge pruning in search performance — a factor which is previously overlooked(Morozov and Babenko, 2018; Liu et al., 2020). Our extensive experiments in §5.2 substantiate this analysis.
@@ 4. Practical Solution
In this section, we introduce a novel $\mathsf{MIPS}$  [^2] For simplicity, we use $\mathsf{MIPS}$ instead of $\epsilon$$\mathsf{MIPS}$ and $k$-$\epsilon$$\mathsf{MIPS}$ for the rest of this paper. framework tailored for practical applications. This framework comprises two key components: a new graph-based indexing structure named Proximity Graph with Spherical Pathways (PSP) and a novel search algorithm with Adaptive Early Termination (AET). The PSP index is specifically designed to address the Biased Solution problem prevalent in $\mathsf{MIPS}$ approaches. Concurrently, the AET mechanism effectively mitigates the Excessive Exploration problem, enhancing search efficiency across diverse query distributions.
@@ 1 $E_{n}\leftarrow\emptyset$; $C\leftarrow$ 2-Hop neighbors of n on $G$; $G^{\prime}\leftarrow G$
Identifying $\mathsf{MIPS}$ neighbors for each node on a large scale is challenging; we address this with a heuristic that involves collecting $k$-hop neighbors and selecting up to $S$ nodes with the largest IP distances as new neighbors. To widen the sensory region, we introduce a smallest-angle constraint inspired by the techniques used in NSSG indexing (Fu et al., 2021). Specifically, any new $\mathsf{MIPS}$ neighbor $m$ of node $n$ (excluding the one with the largest IP distance), the angle $\angle nom\geq\alpha$. Empirically, we find that collecting 2-hop neighbors has significantly enhanced search performance by 8% (§5.2). This process, termed Edge Refinement (EF), is detailed in Algorithm 3.
@@ 4.2. Searching on PSP
1 $R\leftarrow\emptyset$; $Q\leftarrow\emptyset$;
2 $C\leftarrow$ closest navigation cluster look-up;
@@ 9       $f_{p}\leftarrow$ get_features($p$); $\triangleright$ calculate features for $ET(\cdot)$
14      $N_{p}\leftarrow$ neighbors of $p$ in $G$
15       $Q\leftarrow$ batch_insert($Q,N_{p}$)
16       $Q.resize(l_{s})$;$R.resize(k)$triangleright delete small-IP value nodes
17
18
@@ 4.3. Complexity Analysis
The Search complexity for generally favored top-k $\mathsf{MIPS}$ is approximated by $O(c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}+c_{3}f(n)kRn^{1/d}\log n)$ in §3. It is mainly dominated by the $O(kRn^{1/d}\log n)$ term. Notably, our method achieves 95% recall with an average of only 3ms per query on a dataset of 100 million vectors. This efficiency implies a small constant factor in the complexity formula and underscores the practical scalability of our approach for large-scale applications. Empirical evaluations (detailed in §5) aligns with our estimation.
@@ 5.1. Experimental Setup
Datasets. All empirical evaluations were conducted on eight real-world datasets, covering diverse modalities, cardinalities and dimensionalities (Table 2): MNIST (mni, 1998) and Music100 (Morozov and Babenko, 2018; Hu et al., 2008) datasets are the established benchmarks for the $\mathsf{MIPS}$ problem. DBpedia (Oliva and Torralba, 2001) is extracted by OpenAI text-embedding-3-large model. Text2Image (tex, 2021) is a cross-modality dataset where image embeddings (base) are extracted from Se-ResNext-101 and text embeddings (query) from a DSSM Model. Laion10M (Radford et al., 2021b) is a cross-modal retrieval dataset. The embeddings are generated by the CLIP-ViTB/32 model. Commerce100M [^3] Data source will be released upon publication, generated from real traffic logs of a large-scale e-commerce platform. The embeddings are via an advanced method (Fu et al., 2024). The dimension is set to 48 due to the application demands in industrial scenarios, while 16 to 64 are common choices in this field (Zhang et al., 2020; Wang et al., 2021a), discussed in appendix.
@@
Competitors were selected from advanced in-memory methods, encompassing (1) ip-NSW (Morozov and Babenko, 2018): A graph based method using inner-product navigable small world graph. (2) ip-NSW+ (Liu et al., 2020): An enhancement of ip-NSW that introduces an additional angular proximity graph. (3) Möbius-Graph (Zhou et al., 2019): A graph based method that reduces the $\mathsf{MIPS}$ problem to an $\mathsf{NNS}$ problem using Möbius transformation. (4) NAPG (Tan et al., 2021): the state-of-the-art graph based method. (5) Fargo (Zhao et al., 2023a): The latest state-of-the-art LSH based method with theoretical guarantees. (6) ScaNN (Guo et al., 2020): The latest state-of-the-art quantization method, an optimised version based on SOAR (Sun et al., 2024);
Implementation. All baselines except for ScaNN are written in C++. The ScaNN library is called by Python bindings yet written in C++. The experiments are executed on a CentOS machine with 128G RAM on Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz CPU. We use OpenMP for parallel index construction, utilizing 48 threads for all methods. For query execution, we turn off additional optimizations to ensure fair comparison. Our codes are available at (cod, 2025).
@@
5.2. Experimental Results
Exp-1: Theory Verification (Q1). To explore the $\mathsf{MIPS}$-$\mathsf{NNS}$ equivalence with respect to $\mu$, we conduct GNNS on an ideal $\mathsf{PSP}$ index, obtained by setting $L$ to the size of the base vectors and $R$ to infinity in Algorithm 2. This experiment is conducted on MNIST, and we also evaluate Theorem 4 on synthetic datasets sampled from a standard normal distribution, examining the relationship between top-k $\mathsf{MIPS}$ solutions and the top-1 solution’s Euclidean neighbors.
Duality of NNS and MIPS on Proximity Graph. We assessed the average overlap ratio of search paths for $\mathsf{NNS}$ and $\mathsf{MIPS}$, and recall@100 for $\mathsf{MIPS}$, across all queries on MNIST under different $\mu$ (Table 3). The result indicates that a sufficiently large $\mu$ can effectively unify the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paths for all queries. Even sub-optimal $\mu$ yields reasonably good candidates due to highly overlapping paths. Case study in appendix visualizes this. Additionally, we evaluate the overlap ratio between the top-k $\mathsf{MIPS}$ solutions and the neighborhood of the top-1 $\mathsf{MIPS}$ solution. The results in appendix confirm that top-k $\mathsf{MIPS}$ solutions are likely to be distributed in the Euclidean neighborhood of the top-1 $\mathsf{MIPS}$ solution.
@@ 5.2. Experimental Results
Dataset Scale Query Time Indexing Time Index Size
Text2image100K 0.13 (ms) 62 (s) 15 (MB)
Text2image1M 0.46 (ms) 498 (s) 148 (MB)
Text2image10M 0.95 (ms) 5672 (s) 1520 (MB)
@@ 7. Other Related Works
Inner Product (IP) is ubiquitous in metric learning, classification, clustering, knowledge graphs, recommender system, large language model, and information retrieval (Wang et al., 2024; Yu et al., 2014; Fujiwara et al., 2023; Asai et al., 2023; Xu et al., 2020; Huang et al., 2020; Grbovic and Cheng, 2018; Guo et al., 2022; Pan et al., 2023; Zhao et al., 2023b; Asai et al., 2023; Brown et al., 2020; Wang et al., 2021c; Guo et al., 2022). MIPS methods can be categorized by their indexing strategies into LSH, tree, quantization, and graph based approaches:
@@
Tree-based methods. Early tree-based MIPS research (Ram and Gray, 2012; Koenigstein et al., 2012) struggled with high dimensionality. ProMIPS (Song et al., 2021) tackles this issue by projecting the data into lower-dimensional spaces and utilizing iDistance (Jagadish et al., 2005) for indexing. However, it suffers from severe information loss. The recent LRUS-CoverTree (Ma et al., 2024) was designed to mitigate this, while it struggles with solving negative IP values.
Quantization-based methods target to accelerate distance computations with certain approximations. ScanNN (Guo et al., 2020) combine the ”VQ-PQ” framework with an anisotropic quantization loss, presenting a cutting-edge quantization-based solution. Further, SOAR (Sun et al., 2024) uses an orthogonality-amplified residual loss to enhance representations and reaches state-of-the-art and is part of ScaNN library.
Graph-based methods. Proven effective for NNS (Malkov and Yashunin, 2018; Fu et al., 2019, 2021; Lu et al., 2021; Feng et al., 2023; Azizi et al., 2023), graph-based techniques have been adapted for MIPS. ip-NSW (Morozov and Babenko, 2018) replaces Euclidean metric in NSW (Malkov et al., 2014) by IP. ip-NSW+ (Liu et al., 2020) further improves it by adding an angular proximity graph. Möbius-Graph (Zhou et al., 2019) introduces Möbius transformation, albeit with assumptions. Other algorithms  (Tan et al., 2019, 2021) introduce heuristic edge selection strategies to accelerate the search.

### v2

@@ Abstract.
PVLDB, 18(6): 1770 - 1783, 2025.
doi:10.14778/3725688.3725705 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 6 ISSN 2150-8097.
doi:10.14778/3725688.3725705
@@ 1. Introduction
(2) Scalable Solution: We introduce $\mathsf{PSP}$, a scalable graph index with spherical pathways and adaptive search initialization for the biased distribution of $\mathsf{MIPS}$ solutions. We propose an adaptive early termination mechanism to mitigate the excessive exploration.
(3) Extensive Experiments on 8 real-world datasets validate PSP’s theoretical soundness, efficiency, and scalability, highlighting an average 35% speed-up in query process and a 3$\times$ reduction in index size over leading graph methods. PSP is the only one reaching 95% recall under 5ms and thus has been deployed in the large-scale search engines of Shopee, which showcases the application viability.
@@ 2. Preliminaries and Background
While defining $k$-$\mathsf{MIPS}$ and $k$-$\epsilon$$\mathsf{MIPS}$ to find the top $k$ solutions is straightforward, we omit these details here for brevity. Note that the main difference between MIPS and a similar field, Nearest Neighbor Search (NNS), lies in the metric used (Pan et al., 2023) (inner product and Euclidean distance respectively) Recent efforts aim to address both $\epsilon$NNS and $\epsilon$$\mathsf{MIPS}$ problems with similar strategies: reducing the cost of distance calculations and pruning the search space (Zhao et al., 2023a; Ghojogh et al., 2018; Guo et al., 2020; Zheng et al., 2016). To better motivate our study, we revisit the $\epsilon$$\mathsf{MIPS}$ techniques from a novel view: whether transforming the vector space.
@@ 3.1. Equivalence under Scaled Query
We exclude cases where the zero-vector or multiple identical vectors in $\mathcal{D}$ complicate the analysis. Then, we break down the proof step by step.
@@
To align the search paths under both metrics as per Algorithm 1, we can unify the node-selection behavior by localizing the problem: solving for the optimal $p^{*}$ among the neighbors at each greedy step, akin to the conditions established in Theorem 1. Let $l$ be the number of steps in the path. Define $U=\left\{\bar{\mu}_{i}|1\leq i\leq l\right\}$. it’s not hard to verify that there exists a scalar boundary $\sup\left(U\cup\{0\}\right)$ forces the Algorithm 1 under both IP and Euclidean metrics to generate the same search path targeting query $\mu q$. Please refer to (Chen et al., 2025) for the complete proof. ∎
Remarks. From Theorem 1 and Theorem 2, we discern a compelling duality between the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paradigms. These findings suggest that by appropriately scaling the query $q$, we can modulate the overlap in their search paths on a Euclidean proximity graph. By choosing an appropriate $\mu$, the search behavior becomes invariant to the metric used. Our experimental validations further corroborate this theoretical insight on real-world datasets (§3).
@@ 3.2. Efficiency Analyses for Graph-Based MIPS
The complete process of derivation can be found in (Chen et al., 2025). ∎
@@ 7 $G\leftarrow\{G,N\}$
Remarks. Previous transformation-based methods (Shrivastava and Li, 2014, 2015; Neyshabur and Srebro, 2015; Li et al., 2018; Yan et al., 2018; Zhao et al., 2023a; Zhou et al., 2019) sought to reduce the $\mathsf{MIPS}$ problem to the $\mathsf{NNS}$ problem using non-linear transformations, suffering from strong theoretical assumptions (Zhou et al., 2019), structural distortions (Zhao et al., 2023a; Morozov and Babenko, 2018), and complicate data updates (§2), while failed to provide theoretical guarantees on the reachability of the 1-$\mathsf{MIPS}$ solution. Our Theorems 1 and 2 demonstrate that this reachability can be achieved without space transformations. Theorems 3 and 4 validate the efficiency of solving $\mathsf{MIPS}$ on an SSG, highlighting the crucial role of edge pruning in search performance — a factor which is previously overlooked(Morozov and Babenko, 2018; Liu et al., 2020). Our extensive experiments in §3 substantiate this analysis.
@@ 4. Practical Solution
In this section, we introduce a novel $\mathsf{MIPS}$ framework tailored for practical applications. This framework comprises two key components: a new graph-based indexing structure named Proximity Graph with Spherical Pathways (PSP) and a novel search algorithm with Adaptive Early Termination (AET). The PSP index is specifically designed to address the Biased Solution problem prevalent in $\mathsf{MIPS}$ approaches. Concurrently, the AET mechanism effectively mitigates the Excessive Exploration problem, enhancing search efficiency across diverse query distributions.
@@ 5       $S_{m}\leftarrow$ sampling $m_{c}$ nodes from $C[i]$ with Gaussian;
Identifying $\mathsf{MIPS}$ neighbors for each node on a large scale is challenging; we address this with a heuristic that involves collecting $k$-hop neighbors and selecting up to $S$ nodes with the largest IP distances as new neighbors. To widen the sensory region, we introduce a smallest-angle constraint inspired by the techniques used in NSSG indexing (Fu et al., 2021). Specifically, any new $\mathsf{MIPS}$ neighbor $m$ of node $n$ (excluding the one with the largest IP distance), the angle $\angle nom\geq\alpha$. Empirically, we find that collecting 2-hop neighbors has significantly enhanced search performance by 8% (§3). This process, termed Edge Refinement (EF), is detailed in Algorithm 3.
@@ 4.2. Searching on PSP
1 $R\leftarrow\emptyset$; $Q\leftarrow\emptyset$; $C\leftarrow$ closest navigation cluster look-up;
@@ 8       $f_{p}\leftarrow$ get_features($p$); $\triangleright$ calculate features for $ET(\cdot)$
13      $N_{p}\leftarrow$ neighbors of $p$ in $G$; $Q\leftarrow$ batch_insert($Q,N_{p}$);
14       $Q.resize(l_{s})$;$R.resize(k)$; $\triangleright$ delete small-IP value nodes
15
16
@@ 4.3. Complexity Analysis
The Search complexity for generally favored top-k $\mathsf{MIPS}$ is approximated by $O(c_{0}\frac{\log n+c_{1}d}{\log R+c_{2}d}+c_{3}f(n)kRn^{1/d}\log n)$ in §3. It is mainly dominated by the $O(kRn^{1/d}\log n)$ term. Empirical evaluations (detailed in §5) aligns with our estimation.
@@ 5.1. Experimental Setup
Datasets. All empirical evaluations were conducted on eight real-world datasets, covering diverse modalities, cardinalities and dimensionalities (Table 2): MNIST (mni, 1998) and Music100 (Morozov and Babenko, 2018; Hu et al., 2008) datasets are the established benchmarks for the $\mathsf{MIPS}$ problem. DBpedia (Oliva and Torralba, 2001) is extracted by OpenAI text-embedding-3-large model. Text2Image (tex, 2021) is a cross-modality dataset where image embeddings (base) are extracted from Se-ResNext-101 and text embeddings (query) from a DSSM Model. Laion10M (Radford et al., 2021b) is a cross-modal retrieval dataset. The embeddings are generated by the CLIP-ViTB/32 model. Commerce100M, generated from real traffic logs of a large-scale e-commerce platform. The embeddings are via an advanced method (Fu et al., 2024). The dimension is set to 48 due to the application demands in industrial scenarios, while 16 to 64 are common choices in this field (Zhang et al., 2020; Wang et al., 2021a), discussed in (Chen et al., 2025).
Competitors were selected from advanced in-memory methods, encompassing (1) ip-NSW (Morozov and Babenko, 2018): A graph based method using inner-product navigable small world graph. (2) ip-NSW+ (Liu et al., 2020): An enhancement of ip-NSW that introduces an additional angular proximity graph. (3) Möbius-Graph (Zhou et al., 2019): A graph based method that reduces the $\mathsf{MIPS}$ problem to an $\mathsf{NNS}$ problem using Möbius transformation. (4) NAPG (Tan et al., 2021): the state-of-the-art graph based method. (5) Fargo (Zhao et al., 2023a): The latest state-of-the-art LSH based method with theoretical guarantees. (6) ScaNN (Guo et al., 2020): The latest state-of-the-art quantization method, an optimised version based on SOAR (Sun et al., 2024).
Implementation. All baselines except for ScaNN are written in C++. The ScaNN library is called by Python bindings yet written in C++. The experiments are executed on a CentOS machine with 128G RAM on Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz CPU. We use OpenMP for parallel index construction, utilizing 48 threads for all methods. For query execution, we turn off additional optimizations.
@@
Experimental Results Exp-1: Theory Verification (Q1). To explore the $\mathsf{MIPS}$-$\mathsf{NNS}$ equivalence with respect to $\mu$, we conduct GNNS on an ideal $\mathsf{PSP}$ index, obtained by setting $L$ to the size of the base vectors and $R$ to infinity in Algorithm 2. This experiment is conducted on MNIST, and we also evaluate Theorem 4 on synthetic datasets sampled from a standard normal distribution, examining the relationship between top-k $\mathsf{MIPS}$ solutions and the top-1 solution’s Euclidean neighbors.
Duality of NNS and MIPS on Proximity Graph. We assessed the average overlap ratio of search paths for $\mathsf{NNS}$ and $\mathsf{MIPS}$, and recall@100 for $\mathsf{MIPS}$, across all queries on MNIST under different $\mu$ (Table 3). The result indicates that a sufficiently large $\mu$ can effectively unify the $\mathsf{MIPS}$ and $\mathsf{NNS}$ paths for all queries. Even sub-optimal $\mu$ yields reasonably good candidates due to highly overlapping paths. Case study in (Chen et al., 2025) visualizes this. Additionally, we evaluate the overlap ratio between the top-k $\mathsf{MIPS}$ solutions and the neighborhood of the top-1 $\mathsf{MIPS}$ solution. The results in (Chen et al., 2025) confirm that top-k $\mathsf{MIPS}$ solutions are likely to be distributed in the Euclidean neighborhood of the top-1 $\mathsf{MIPS}$ solution.
@@
Dataset Scale Query Time Indexing Time Index Size
Text2image100K 0.13 (ms) 62 (s) 15 (MB)
Text2image1M 0.46 (ms) 498 (s) 148 (MB)
Text2image10M 0.95 (ms) 5672 (s) 1520 (MB)
@@ 7. Other Related Works
Inner Product (IP) is ubiquitous in metric learning, classification, clustering, knowledge graphs, recommender system, large language model, and information retrieval (Yu et al., 2014; Fujiwara et al., 2023; Xu et al., 2020; Huang et al., 2020; Grbovic and Cheng, 2018; Pan et al., 2023; Zhao et al., 2023b; Wang et al., 2021c; Guo et al., 2022; Wang et al., 2024) MIPS methods can be categorized by their indexing strategies into LSH, tree, quantization, and graph based approaches:
@@
Tree-based methods. Early tree-based MIPS research (Ram and Gray, 2012; Koenigstein et al., 2012) struggled with high dimensionality. ProMIPS (Song et al., 2021) tackles this issue by projecting the data into lower-dimensional spaces but it suffers from severe information loss. LRUS-CoverTree (Ma et al., 2024) was designed to mitigate this, while it struggles with solving negative IP values.
Quantization-based methods target to accelerate distance computations with certain approximations. ScanNN (Guo et al., 2020) combine the ”VQ-PQ” framework with an anisotropic quantization loss, presenting a cutting-edge quantization-based solution. SOAR (Sun et al., 2024) uses an orthogonality-amplified residual loss to enhance representations and reaches state-of-the-art and is part of ScaNN library.
Graph-based methods. Proven effective for NNS (Malkov and Yashunin, 2018; Fu et al., 2019, 2021; Lu et al., 2021; Feng et al., 2023; Azizi et al., 2023), graph-based techniques have been adapted for MIPS. ip-NSW (Morozov and Babenko, 2018) replaces Euclidean metric in NSW (Malkov et al., 2014) by IP, and ip-NSW+ (Liu et al., 2020) further improves it by adding an angular proximity graph. Möbius-Graph (Zhou et al., 2019) introduces Möbius transformation, albeit with assumptions. Other algorithms  (Tan et al., 2019, 2021) introduce heuristic edge selection strategies to accelerate the search.